# Emulation of Complex Matrix Multiplication based on the Chinese Remainder Theorem

### v1

@@ I-A Background
General-purpose computing architectures have increasingly adopted low-precision matrix multiplication units in Graphics Processing Units (GPUs), Tensor Processing Units (TPUs), and Neural Processing Units (NPUs). Representative examples include Tensor Cores in NVIDIA GPUs [tensorcore], Matrix Cores in AMD GPUs [matrixcore], the MXU in Google TPUs [tpu], and the MAC array in Intel NPUs [intelXecore]. These units offer substantially higher arithmetic throughput than high-precision units and have become central components of modern accelerator design. Table I shows the theoretical throughput performance in TFLOPS or TOPS of AMD [MI300A, MI300X, MI325X, MI350X, MI355X] and NVIDIA GPUs [A100, H100, Blackwell, RTX5080]. In recent years, such units have also begun to appear as standard features even in laptop-class processors [copilot+pc, SME], further broadening their availability. While these units were originally motivated by machine-learning workloads, their architectural characteristics also present an opportunity to accelerate high-precision computations by emulating them on low-precision hardware. Consequently, emulation-based high-precision matrix multiplication has emerged as an active research topic in the high-performance computing community [brower2025mixedprecisionabinitiotensor, abdelfattah2025analysisfloatingpointmatrixmultiplication, schwarz2025guaranteed, mukunoki2025sparseiterativesolversusing, dongarra2024hardwaretrendsimpactingfloatingpoint, 11196192, 11196413].
TABLE I: TFLOPS and TOPS of AMD and NVIDIA GPUs [MI300A, MI300X, MI325X, MI350X, MI355X, A100, H100, Blackwell, RTX5080]
@@ I-B Contributions
A prior line of work introduced the Ozaki-I [ozaki2012error, ozaki2013generalization, mukunoki2020, mukunoki2025dgemmfp64arithmetic, ootomo2024dgemm, uchino2025Performance] and Ozaki-II [ozaki-scheme2, uchino_ozaki2] schemes as frameworks for emulating high-precision matrix multiplication using low-precision matrix engines, demonstrating high performance and power efficiency on contemporary GPUs. In parallel, the Ozaki-I scheme has recently been incorporated into cuBLAS (from version 13.0 Update 2), and this implementation supports complex matrix multiplication. However, for real-valued matrix multiplication, previous studies have reported that the Ozaki-II scheme achieves higher performance than Ozaki-I. This suggests that a complex extension of the Ozaki-II scheme would likewise outperform the complex Ozaki-I approach. Despite this expectation, no complex-valued formulation of Ozaki-II has been developed to date. This absence is notable because complex matrix multiplication is a key computational kernel in many HPC applications, and its performance can dominate the overall runtime of large-scale scientific simulations.
@@
To address this gap, the present work extends the Ozaki-II scheme to single- and double-precision complex matrix multiplication (CGEMM and ZGEMM, respectively). We develop emulation algorithms for CGEMM and ZGEMM by combining the Ozaki-II scheme with a Karatsuba-based formulation of complex arithmetic. The resulting methods maintain numerical accuracy comparable to standard CGEMM and ZGEMM while effectively exploiting the high throughput of low-precision matrix engines. In addition, several improvements applicable to existing SGEMM and DGEMM emulation techniques are introduced, further enhancing the efficiency of the Ozaki schemes in the real domain. The proposed algorithms are released as a portable library supporting both CUDA and HIP backends, enabling execution on NVIDIA GPUs and AMD GPUs. Performance is evaluated on an NVIDIA GH200 Grace Hopper Superchip, an NVIDIA B200 Blackwell GPU, an NVIDIA GeForce RTX 5080 Blackwell GPU, and an AMD Instinct MI300X GPU, and compared against cuBLAS and hipBLAS implementations. For sufficiently large matrix sizes, the proposed methods surpass both vendor-optimized libraries and, in particular, achieve higher throughput than the cuBLAS Ozaki-I–based ZGEMM emulation, demonstrating the performance potential of Ozaki-II-based emulation for CGEMM and ZGEMM. We also provide performance models for the proposed methods, and the predicted throughput closely aligns with the measured results across all tested configurations.
@@
The remainder of this paper is organized as follows. Section II introduces the Ozaki-II scheme for real-valued matrix multiplication. Section III presents the proposed efficient methods for emulating CGEMM and ZGEMM. Section IV shows the numerical results comparing the proposed emulation with the vender-optimized GEMM routines. Section V presents the concluding remarks. Note that in this paper, we do not consider matrix shapes for which performance becomes memory-bound, such as small matrices or highly skinny matrices, because low-precision matrix engines do not achieve high throughput on such workloads.
@@ II Ozaki-II Scheme for Real Matrix Multiplication
The Ozaki-II scheme is an emulation method of matrix multiplication based on a large-integer multiplication technique using the Chinese Remainder Theorem (CRT). The scheme for matrix multiplication of $A\in\mathbb{R}^{m\times k}$ and $B\in\mathbb{R}^{k\times n}$ consists of the following three steps:
@@
    Convert $C^{\prime}$ back to floating-point format $C\in\mathbb{R}^{m\times n}$ by inversely scaling.
The integer matrix multiplication in step 2 is performed by decomposing the elements of $A^{\prime}$ and $B^{\prime}$ into residues with respect to a predefined set of pairwise coprime moduli. Let $p_{1},p_{2},\dots,p_{N}\in\mathbb{N}$ denote these moduli and define $P:=\prod_{\ell=1}^{N}p_{\ell}$. Let $\mathrm{mod}(x,p_{\ell})$ for $x\in\mathbb{Z}$ denote the symmetric remainder of $x/p_{\ell}$, i.e., $\mathrm{mod}(x,p_{\ell})=x-p_{\ell}\cdot\mathrm{round}(x/p_{\ell})$, where $\mathrm{round}(\cdot)$ rounds its input to the nearest integer. Each element of $A^{\prime}$ and $B^{\prime}$ is mapped to its residue representation modulo each $p_{\ell}$ as
@@
is satisfied for all $(i,j)$ pairs, the possible value of $c^{\prime}_{ij}$ is only $\mathrm{mod}\left((A^{\prime}B^{\prime})_{ij},P\right)$ and the CRT guarantees that the reconstructed $C^{\prime}$ is unique. Therefore, the scaling vector used in step 1 must be chosen such that the bound (4) is ensured for $A^{\prime}$ and $B^{\prime}$. Increasing the number of moduli $p_{\ell}$ enlarges $P$, thereby relaxing the uniqueness condition. However, this also increases the number of matrix multiplications required, leading to higher computational cost.
In previous work [uchino_ozaki2], an algorithm was proposed in which each modular matrix multiplication (3) is executed using INT8 matrix engines. Algorithm 1 shows the outline. Hereafter, we assume that $k\leq 2^{17}$ to perform INT8 matrix multiplication without error. In [uchino_ozaki2], step iv at the line 7 of Algorithm 1 uses the standard (non-symmetric) modulo operation that returns values in $0,\,p_{\ell})$ and $E^{\prime}_{\ell}$ are UINT8 matrices. In DGEMM emulation, step v at the line 7 evaluates $S=\sum_{\ell=1}^{N}P/p_{\ell}\cdot q_{\ell}\cdot E^{\prime}_{\ell}$ by the unevaluated sum $S_{1}+S_{2}$ as
@@
where $s_{\ell 1}$ contains the upper $(53-8-\lceil\log_{2}N\rceil)$ bits of $P/p_{\ell}\cdot q_{\ell}$, $s_{\ell 2}$ stores the remaining lower bits rounded to double precision, and $s_{\ell 1}+s_{\ell 2}\approx P/p_{\ell}\cdot q_{\ell}$. In (5), the value of $s_{\ell 1}$ is limited by the precision of the double-precision significand. Since each element of $E^{\prime}_{\ell}$ is an 8-bit unsigned integer, the product $s_{\ell 1}E^{\prime}_{\ell}$ requires 8 bits of significand to represent the contribution of $E^{\prime}_{\ell}$. Moreover, accumulating $N$ such terms requires an additional $\lceil\log_{2}N\rceil$ bits to avoid overlap between partial sums. Thus, the high-order part $s_{\ell 1}$ may occupy at most $(53-8-\lceil\log_{2}N\rceil)$ bits of precision. This bit allocation ensures that all products $s_{\ell 1}E^{\prime}_{\ell}$ can be accumulated in $S_{1}$ without rounding error. In contrast, the present work adopts a symmetric modulo operation in step iv at the line 7, ensuring that each $E^{\prime}_{\ell}$ becomes INT8 from UINT8. Consequently, the upper part $s_{\ell 1}$ of $P/p_{\ell}\cdot q_{\ell}$ can use $(53-7-\lceil\log_{2}N\rceil)$ bits instead of $(53-8-\lceil\log_{2}N\rceil)$ bits, gaining one additional effective bit compared with the previous method. This slight but meaningful increase in representable precision improves the numerical robustness of the reconstruction step without changing the overall algorithmic structure.
@@
[…]
3:Determine the moduli $p_{\ell}\leq 256$ for $\ell=1,\dots,N$.
4:Determine modular multiplicative inverses $q_{\ell}$ of $P/p_{\ell}$ modulo $p_{\ell}$ for $\ell=1,\dots,N$.
5:Determine the scaling vectors $\mu\in\mathbb{R}^{m}$ and $\nu\in\mathbb{R}^{n}$.
6:Convert AA and $B$ to $A^{\prime}\in\mathbb{Z}^{m\times k}$ and $B^{\prime}\in\mathbb{Z}^{k\times n}$:
@@
7: Compute $C^{\prime}\approx A^{\prime}B^{\prime}$ based on the CRT:
@@
8:Convert $C^{\prime}$ back to floating-point format $C\in\mathbb{R}^{m\times n}$:
@@
To satisfy the condition (4) efficiently, the scaling vectors used in step 1 must be determined appropriately. Previous work introduced two strategies for computing this scaling: a fast mode and an accurate mode. The fast mode estimates an upper bound of $\sum_{h=1}^{k}|a^{\prime}_{ih}||b^{\prime}_{hj}|$ via the Cauchy–Schwarz inequality and determines the scaling vector analytically based on this estimate. This approach incurs minimal preprocessing cost, but the bound obtained from the Cauchy–Schwarz inequality tends to be overly conservative. As a consequence, a larger $P$ (i.e., more moduli) is often required to satisfy the uniqueness condition, increasing the computational workload in step 2. In contrast, the accurate mode estimates the upper bound by performing an auxiliary matrix multiplication using INT8 matrix engines. Although this mode requires additional computation during step 1, the direct estimation yields a much tighter bound than that obtained via the Cauchy–Schwarz inequality. This typically enables the use of fewer moduli, reducing the number of modular matrix multiplications and potentially improving overall performance.
@@ III Proposed Methods
This section presents the extension of the Ozaki-II scheme to complex matrix multiplication, targeting high-performance execution on INT8 matrix engines. Let $A=A_{R}+iA_{I}\in\mathbb{C}^{m\times k}$ for $A_{R},A_{I}\in\mathbb{R}^{m\times k}$ and $B=B_{R}+iB_{I}\in\mathbb{C}^{k\times n}$ for $B_{R},B_{I}\in\mathbb{R}^{k\times n}$, and we compute $C=C_{R}+iC_{I}:=AB$, where $C_{R},C_{I}\in\mathbb{R}^{m\times n}$. A standard approach to handle complex matrix multiplication is to employ the following representation:
@@ III-A INT8 Complex Matrix Multiplication
The formulations in (7) and (8) allow complex matrix multiplication to be computed using a single real GEMM applied to an expanded matrix. For the formulations, the applicability condition on the inner dimension changes from $k\leq 2^{17}$ to $k\leq 2^{16}$. Another standard approach to handle complex matrix multiplication is the Karatsuba algorithm. The Karatsuba algorithm expresses the complex product using three real matrix multiplications as
@@
Figure 1: Performance comparison of four INT8-based GEMM strategies on NVIDIA H100 NVL with CUDA Toolkit 12.8.61. The first two methods use a single INT8 matrix multiplication with dimensions $(2h,h,2h)$ and $(h,2h,2h)$, respectively. The third method performs three INT8 multiplications of size $(h,h,h)$, while the fourth applies the same three-multiplication scheme but introduces blocking along the $n$ dimension.

### v3

@@ I-A Background
General-purpose computing architectures have increasingly adopted low-precision matrix multiplication units in graphics processing units (GPUs), tensor processing units (TPUs), and neural processing units (NPUs). Representative examples include Tensor Cores in NVIDIA GPUs [tensorcore], Matrix Cores in AMD GPUs [matrixcore], Matrix Multiply Units in Google TPUs [tpu], and Multiply-Accumulate arrays in Intel NPUs [intelXecore]. These units offer substantially higher arithmetic throughput than that of high-precision units and have become central components of modern accelerator design. Table I shows the theoretical throughput performance for AMD [MI300A, MI300X, MI325X, MI350X, MI355X] and NVIDIA GPUs [A100, H100, Blackwell, RTX5080]. In recent years, low-precision matrix multiplication units have also begun to appear as standard features even in laptop-class processors [copilot+pc, SME], further broadening their availability. While these units were originally motivated by machine-learning workloads, their architectural characteristics also present an opportunity to accelerate high-precision computations by emulating such computations on low-precision hardware. Consequently, emulation-based high-precision matrix multiplication has emerged as an active research topic in the high-performance computing community [brower2025mixedprecisionabinitiotensor, abdelfattah2025analysisfloatingpointmatrixmultiplication, schwarz2025guaranteed, mukunoki2025sparseiterativesolversusing, dongarra2024hardwaretrendsimpactingfloatingpoint, 11196192, 11196413].
TABLE I: Performance (in TFLOPS or TOPS where appropriate) of AMD and NVIDIA GPUs [MI300A, MI300X, MI325X, MI350X, MI355X, A100, H100, Blackwell, RTX5080]
@@ I-B Contributions
A prior line of work introduced the Ozaki-I [ozaki2012error, ozaki2013generalization, mukunoki2020, mukunoki2025dgemmfp64arithmetic, ootomo2024dgemm, uchino2025Performance] and Ozaki-II [ozaki-scheme2, uchino_ozaki2] schemes as frameworks for emulating high-precision matrix multiplication using low-precision matrix engines, demonstrating high performance and power efficiency on contemporary GPUs. The Ozaki-I scheme has recently been incorporated into cuBLAS (from version 13.0 Update 2). This implementation supports complex matrix multiplication. However, for real-valued matrix multiplication, previous studies have reported that the Ozaki-II scheme achieves higher performance than Ozaki-I. This suggests that a complex extension of the Ozaki-II scheme would likewise outperform the complex Ozaki-I approach. Despite this expectation, no complex-valued formulation of the Ozaki-II scheme has been developed. This absence is notable because complex matrix multiplication is a key computational kernel in many high-performance computing applications and its performance can dominate the overall runtime of large-scale scientific simulations.
@@
To address this gap, the present work extends the Ozaki-II scheme to single- and double-precision complex matrix multiplication (CGEMM and ZGEMM, respectively). We develop emulation algorithms for CGEMM and ZGEMM by combining the Ozaki-II scheme with a Karatsuba-based formulation of complex arithmetic. The developed methods maintain numerical accuracy comparable to that of standard CGEMM and ZGEMM while effectively exploiting the high throughput of low-precision matrix engines. In addition, several improvements applicable to existing single- and double-precision general matrix multiplication (GEMM) emulation techniques are introduced, further enhancing the efficiency of the Ozaki schemes in the real domain. The proposed algorithms are released as a portable library that supports both CUDA and HIP backends, enabling execution on NVIDIA and AMD GPUs. Performance is evaluated on an NVIDIA GH200 Grace Hopper Superchip, an NVIDIA B200 Blackwell GPU, an NVIDIA GeForce RTX 5080 Blackwell GPU, and an AMD Instinct MI300X GPU and compared against cuBLAS and hipBLAS implementations. For sufficiently large matrix sizes, the proposed methods surpass both vendor-optimized libraries and, in particular, achieve higher throughput than that of cuBLAS Ozaki-I-based ZGEMM emulation, demonstrating the performance potential of Ozaki-II-based emulation for CGEMM and ZGEMM. We also provide performance models for the proposed methods. The predicted throughput closely aligns with the measurement results across all tested configurations.
@@
The remainder of this paper is organized as follows. Section II introduces the Ozaki-II scheme for real-valued matrix multiplication. Section III presents the proposed efficient methods for emulating CGEMM and ZGEMM. Section IV shows numerical results comparing the proposed emulation with vender-optimized GEMM routines. Section V presents the concluding remarks. Note that in this paper, we do not consider matrix shapes for which performance becomes memory-bound, such as small matrices or very skinny matrices, because low-precision matrix engines do not achieve high throughput on such workloads.
@@ II Ozaki-II Scheme for Real Matrix Multiplication
The Ozaki-II scheme is an emulation method for matrix multiplication based on a large-integer multiplication technique using the Chinese remainder theorem (CRT). The scheme for the multiplication of matrices $A\in\mathbb{R}^{m\times k}$ and $B\in\mathbb{R}^{k\times n}$ consists of the following three steps:
@@
    Convert $C^{\prime}$ back to floating-point format $C\in\mathbb{R}^{m\times n}$ via inverse scaling.
The integer matrix multiplication in step 2 is performed by decomposing the elements of $A^{\prime}$ and $B^{\prime}$ into residues with respect to a predefined set of pairwise coprime moduli. Let $p_{1},p_{2},\dots,p_{N}\in\mathbb{N}$ denote these moduli and define $P:=\prod_{\ell=1}^{N}p_{\ell}$. Let $\mathrm{mod}(x,p_{\ell})$ for $x\in\mathbb{Z}$ denote the symmetric remainder of $x/p_{\ell}$ (i.e., $\mathrm{mod}(x,p_{\ell})=x-p_{\ell}\cdot\mathrm{round}(x/p_{\ell})$, where $\mathrm{round}(\cdot)$ rounds its input to the nearest integer). Each element of $A^{\prime}$ and $B^{\prime}$ is mapped to its residue representation modulo each $p_{\ell}$ as
@@
is satisfied for all $(i,j)$ pairs, the only possible value of $c^{\prime}_{ij}$ is $\mathrm{mod}\left((A^{\prime}B^{\prime})_{ij},P\right)$ and the CRT guarantees that the reconstructed $C^{\prime}$ is unique. Therefore, the scaling vector used in step 1 must be chosen such that the bound (4) is ensured for $A^{\prime}$ and $B^{\prime}$. Increasing the number of moduli $p_{\ell}$ enlarges $P$, thereby relaxing the uniqueness condition. However, this also increases the number of matrix multiplications required, leading to higher computational cost.
In previous work [uchino_ozaki2], an algorithm was proposed in which each modular matrix multiplication (3) is executed using INT8 matrix engines. Algorithm 1 shows the outline. Hereafter, we assume that $k\leq 2^{17}$ to perform INT8 matrix multiplication without error. In [uchino_ozaki2], step V-iv of Algorithm 1 uses the standard (non-symmetric) modulo operation that returns values in $0,\,p_{\ell})$ and $E^{\prime}_{\ell}$ are UINT8 matrices. In DGEMM emulation, step V-v of Algorithm 1 evaluates $S=\sum_{\ell=1}^{N}P/p_{\ell}\cdot q_{\ell}\cdot E^{\prime}_{\ell}$ using the unevaluated sum $S_{1}+S_{2}$ as
@@
where $s_{\ell 1}$ contains the upper $(53-8-\lceil\log_{2}N\rceil)$ bits of $P/p_{\ell}\cdot q_{\ell}$, $s_{\ell 2}$ stores the remaining lower bits rounded to double precision, and $s_{\ell 1}+s_{\ell 2}\approx P/p_{\ell}\cdot q_{\ell}$. In (5), the value of $s_{\ell 1}$ is limited by the precision of the double-precision significand. Since each element of $E^{\prime}_{\ell}$ is an 8-bit unsigned integer, the product $s_{\ell 1}E^{\prime}_{\ell}$ requires 8 bits of significand to represent the contribution of $E^{\prime}_{\ell}$. Moreover, accumulating $N$ such terms requires an additional $\lceil\log_{2}N\rceil$ bits to avoid overlap between partial sums. Thus, the high-order part $s_{\ell 1}$ may occupy at most $(53-8-\lceil\log_{2}N\rceil)$ bits of precision. This bit allocation ensures that all products $s_{\ell 1}E^{\prime}_{\ell}$ can be accumulated in $S_{1}$ without rounding error. In contrast, the present work adopts a symmetric modulo operation in step V-iv of Algorithm 1, ensuring that each $E^{\prime}_{\ell}$ becomes INT8 from UINT8. Consequently, the upper part $s_{\ell 1}$ of $P/p_{\ell}\cdot q_{\ell}$ can use $(53-7-\lceil\log_{2}N\rceil)$ bits instead of $(53-8-\lceil\log_{2}N\rceil)$ bits, gaining one additional effective bit compared to the previous method. This slight but meaningful increase in representable precision improves the numerical robustness of the reconstruction step without changing the overall algorithmic structure.
@@
$A\in\mathbb{R}^{m\times k}$, $B\in\mathbb{R}^{k\times n}$, $N\in\mathbb{N}$
$C\approx AB\in\mathbb{R}^{m\times n}$
Step I
Determine the moduli $p_{\ell}\leq 256$ for $\ell=1,\dots,N$.
Step II
Determine modular multiplicative inverses $q_{\ell}$ of $P/p_{\ell}$ modulo $p_{\ell}$ for $\ell=1,\dots,N$.
Step III
Determine the scaling vectors $\mu\in\mathbb{R}^{m}$ and $\nu\in\mathbb{R}^{n}$.
Step IV
Convert $A$ and $B$ to $A^{\prime}\in\mathbb{Z}^{m\times k}$ and $B^{\prime}\in\mathbb{Z}^{k\times n}$:
@@
Step V
Compute $C^{\prime}\approx A^{\prime}B^{\prime}$ based on the CRT:
@@
Step VI
Convert $C^{\prime}$ back to floating-point format $C\in\mathbb{R}^{m\times n}$:
@@
To satisfy the condition (4) efficiently, the scaling vectors used in step IV of Algorithm 1 must be determined appropriately. Previous work introduced two strategies for computing this scaling: a fast mode and an accurate mode. The fast mode estimates the upper bound of $\sum_{h=1}^{k}|a^{\prime}_{ih}||b^{\prime}_{hj}|$ via the Cauchy–Schwarz inequality and determines the scaling vector analytically based on this estimate. This approach incurs minimal preprocessing cost, but the obtained bound tends to be overly conservative. As a consequence, a larger $P$ (i.e., more moduli) is often required to satisfy the uniqueness condition, increasing the computational workload in step V Algorithm 1. In contrast, the accurate mode estimates the upper bound by performing an auxiliary matrix multiplication using INT8 matrix engines. Although this mode requires additional computation during step III of Algorithm 1, the direct estimation yields a much tighter bound than that obtained via the Cauchy–Schwarz inequality. This typically enables the use of fewer moduli, reducing the number of modular matrix multiplications and potentially improving overall performance.
@@ III Proposed Methods
This section presents the proposed extension of the Ozaki-II scheme to complex matrix multiplication, targeting high-performance execution on INT8 matrix engines. Let $A=A_{R}+iA_{I}\in\mathbb{C}^{m\times k}$ for $A_{R},A_{I}\in\mathbb{R}^{m\times k}$ and $B=B_{R}+iB_{I}\in\mathbb{C}^{k\times n}$ for $B_{R},B_{I}\in\mathbb{R}^{k\times n}$. We compute $C=C_{R}+iC_{I}:=AB$, where $C_{R},C_{I}\in\mathbb{R}^{m\times n}$. A standard approach for handling complex matrix multiplication is to employ the following representation:
@@ III-A INT8 Complex Matrix Multiplication
The formulations in (7) and (8) allow complex matrix multiplication to be computed using a single real GEMM applied to an expanded matrix. For the formulations, the applicability condition on the inner dimension changes from $k\leq 2^{17}$ to $k\leq 2^{16}$. Another standard approach for handling complex matrix multiplication is the Karatsuba algorithm. The Karatsuba algorithm expresses the complex product using three real matrix multiplications as
@@
Figure 1: Performance comparison of four INT8-based GEMM strategies on NVIDIA H100 NVL (CUDA Toolkit 12.8.61). The first two methods use a single INT8 matrix multiplication with dimensions $(2h,h,2h)$ and $(h,2h,2h)$, respectively. The third method performs three INT8 multiplications of size $(h,h,h)$. The fourth method applies the same three-multiplication scheme but introduces blocking along the $n$ dimension.